[Hallc_running] [New Logentry] Status Update

pooser at jlab.org pooser at jlab.org
Sat Feb 24 16:10:01 EST 2018

Logentry Text:
Owl shifts for the near future have been filled, muchos kudos to Dave M!

During the night we remained on LH2 with the SHMS at -2.5, 39 deg and HMS at -1.5 and 59 deg.  At ~8 am we switched to LD2 and at ~3 pm we switched to 10 cm dummy.  Beam since last night has been relatively stable (https://logbooks.jlab.org/entry/3537839) with no sizable down time periods encountered. For the current kinematic setting we have collected ~270k e-’s (of 550k) in the SHMS and ~10k e-’s (of 25k) in the HMS with a LD2 target, ~245k e-’s (of 550k) in the SHMS and ~6.1k e-’s (of 25k) in the HMS with the LH2 Target.  For the 10 cm dummy target we have collected ~18k e-’s (of 55k) in the SHMS and ~0.8k e-’s (of 3k) in the HMS.  For a full synopsis of the data taken heretofore follow the following link: https://docs.google.com/document/d/13i-SQGGES52d8fjJo9CTdgMc0A7J7z_t9FdKewW6HEM/edit?ts=5a8db865.

The plan moving forward is to complete the run plan (https://hallcweb.jlab.org/DocDB/0009/000953/001/F2_EMC_runlist_final.pdf) at the current kinematic setting for the SHMS while the HMS will remain in the same configuration for the foreseeable future.  As a very rough estimate for the current SHMS setting, we require an approximate 10 hrs on LD2, 14 hrs on LH2, and 5 hrs on dummy prior to changing the momentum setting to the next momentum setting at -2.0 GeV.  If all goes well we should be changing SHMS momentum settings some time close to beginning of day shift Monday morning.

MCC plans to conduct beam studies beginning at 9 am on Monday morning.  During the first ~2 hours of the beam studies MCC will perform fast feedback work which results in no beam and no access into Hall C.  During the following ~2 hours Hall C can receive beam parasitically which we will happily take.  For the remainder of the beam studies optics measurements will be conducted there will be no parasitic beam available for Hall-C however, we will utilize this time in order to make the required accesses into the Hall.  If you have any work that needs to be conducted inside the Hall which that can be completed inside of 4 hours please let me know as soon as possible so that I may plan accordingly.

Work currently planned during Monday’s (2/26) beam studies starting at 9 am:
Joe Beaufait: diagnose (and potentially fix) HMS door/shutter breaker (https://logbooks.jlab.org/entry/3535292)
Andrew: install five Y strainers on the SHMS power supply LCW lines.  The power supplies will have to be ramped down and turned off

Special notices in effect for shift crews:
Trouble shooting HMS Q2 power supply: https://logbooks.jlab.org/entry/3537306
Don't ramp HMS Q2 higher than 610A: https://logbooks.jlab.org/entry/3535714
TO Instructions for LH2 <--> LD2 switches : https://logbooks.jlab.org/entry/3534487
Maximum Boron targets is 40 muA: https://logbooks.jlab.org/entry/3535297
Any missing reference time messages should not be ignored: https://logbooks.jlab.org/entry/3535289

Recent analysis work:
The BCM gains have been update in hallc-online/hallc_replay: https://logbooks.jlab.org/entry/3537764 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3537872
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180224/0ea92a34/attachment.html>

More information about the Hallc_running mailing list